A driver takes 0.20 s to apply the brakes after he sees a need for it. This is called the reaction time of the driver. If he is driving a car at a speed of 54 km/h and the brakes cause a deceleration of 6.0 m/s2, find the distance travelled by the car after he see the need to put the brakes on.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. In reaction time the body moves with the speed 54 km/hr = 15 m/sec (constant speed) Distance travelled in this time is S1 = 15 × 0.2 = 3 m. When brakes are applied, u = 15 m/s, v = 0, a = –6 m/s2 (deceleration) S2 = (V^2-u^2)/2a = (0-〖15〗^2)/(2(-6)) = 18.75 m Total distance s = s1 + s2 = 3 + 18.75 = 21.75 = 22 m.